CAVITY RINGDOWN SPECTROSCOPY ON THE b^{1}Σ_{g}^{+} - X^{3}Σ_{g}^{−} (v’, 0) BANDS (v’=0-3) OF O_{2}-ISOTOPOMERS

\underline{HANS NAUS} and WIM UBACHS, Vrije Universiteit, Department of Physics and Astronomy, De Boelelaan 1081, 1081 HV Amsterdam, The Netherlands.

Cavity RingDown Spectroscopy (CRDS) has been used to investigate the b^{1}Σ_{g}^{+} - X^{3}Σ_{g}^{−} system of oxygen isotopomers. The absorption bands of this system are very weak due to the strongly forbidden character of gerade-gerade and Σ^{+} − Σ^{−} transitions, and they can only be observed via a magnetic dipole transition moment. The four allowed branches are a factor of 10^{9} weaker than typical electric dipole allowed transitions. The high sensitivity (10^{−9} cm^{−1}), obtained even in a short cell (< 1 m), makes CRDS ideal for spectroscopy on isotopically enriched samples. Our studies yield updated or new values for the molecular constants of the b^{1}Σ_{g}^{+} - X^{3}Σ_{g}^{−} (v’, 0) bands for v’= 0-3 of oxygen isotopomers. These values may help to identify emission bands from excited vibrational levels of O_{2} (b^{1}Σ_{g}^{+}), including minority isotopomers, as observed in the Earth’s atmosphere^{a} and can aid in the interpretation of weak signals in the search for possible violation of the symmetrization postulate in ^{16}O_{2}^{b}^{c}.

aT. G. Slanger, D. L. Huestis, D. E. Osterbrock and J. P. Fullbright Science \underline{277}, 1485, 1997.

bM. de Angelis, G. Gagliardi, L. Gianfrani and G. M. Tino Phys. Rev. Lett. \underline{76}, 2840, 1996.

cR. C. Hilborn and C. L. Yuca Phys. Rev. Lett. \underline{76}, 2844, 1996.